OpenStudy (alexandervonhumboldt2):

The volume of the main tank is 1436026.6666666666666666666666667 feet³. If the maximum density of killer whales per cubic foot is 0.000011142, what is the maximum number of killer whales allowed in the main show tank at any given time?

OpenStudy (alexandervonhumboldt2):

@chosenmatt 2jim_thompson5910 @Mertsj @perl @SolomonZelman

OpenStudy (anonymous):

WE can figure out the mass of killer whales, but not necessarily the quantity

OpenStudy (alexandervonhumboldt2):

ok so how?

OpenStudy (perl):

density = (# killer whales) / volume

OpenStudy (mertsj):

\[1426026\frac{2}{3} ft^3 \times\frac{1.1142 \times 10^{-5}killer whales}{1 ft^3}\]

OpenStudy (anonymous):

OH, did they change density to whales/volume? lol

OpenStudy (alexandervonhumboldt2):

no i think

OpenStudy (alexandervonhumboldt2):

the same

OpenStudy (alexandervonhumboldt2):

thats not relaistic but...

OpenStudy (mertsj):

yes. Sorry for the typo...old age, old eyes

OpenStudy (alexandervonhumboldt2):

@Mertsj so thats the formula by which i can count right?

OpenStudy (anonymous):

Use @perl equation... multiply both sides by volume.

OpenStudy (mertsj):

It's just a unit conversion problem. Convert a certain number of cubic feet to whales because you know the number of whales per cubic foot.

OpenStudy (alexandervonhumboldt2):

so 0.000011142=#whales/436026.6666666666666666666666667 right?

OpenStudy (perl):